# BLaST: High Performance Inference and Pretraining using BLock Sparse Transformers

### v1

@@ 5.2. BLaST for Inference
The GLUE benchmark assesses the general language understanding capabilities of models. Specifically, we evaluate on five tasks—Recognizing Textual Entailment (RTE), the Multi-Genre Natural Language Inference Corpus (MNLI), The Stanford Sentiment Treebank (SST-2), the Winograd Schema Challenge (WNLI), and the Corpus of Linguistic Acceptability (CoLA)—and report the average score calculated as the arithmetic mean of all other task specific metrics. We observe that BLaST is robust to variations in block size, achieving a similar average score across the five tasks regardless of the chosen block size. Furthermore, BLaST maintains an accuracy comparable to that of the dense baseline. For instance, on the WNLI dataset, BLaST achieves performance on par with the dense model across all settings, and on the CoLA dataset, the variants BLaST-70%/32x32, BLaST-80%/32x32, and BLaST-70%/128x128 outperform the dense baseline.
@@
Collectively, these results demonstrate the robustness of BLaST to varying sparsity levels, and block sizes for reconstructing accuracy, making BLaST a ready to use method for pretrained networks when speedup and memory reduction are of importance.
5.2.1. Memory Footprint for Inference
@@
The pruning of the weights by up to 95% can have a large impact on the total number of parameters of the model, and with that affect model’s memory footprint. Figure 7 shows the number of NVIDIA GH200 GPUs that are necessary for storing the model weights, assuming 96GB HBM per GPU and FP32 weights. BLaST is capable of reducing the number of necessary GPUs to almost a third of the dense Llama 3.2 405B. This directly translates into reduced capital investments required for establishing inference data centers and lower energy consumption for their operation.
@@
Figure 7. Number of GH200 GPUs required for storing weights in FP32 of the given model assuming 96GB per GPU.
@@ 5.3. BLaST for Pretraining
(a) GPT2-XL pretraining with 4.9B tokens.
(b) Llama 3.2 pretraining with 4.2B tokens.
Figure 8. Time per iteration for GPT2-XL and Llama 3.2 with OpenWebText (Gokaslan2019, ). Each line plot uses the hyper parameters from Table 2.
5.3.1. Pretraining LLMs
We pretrain GPT2 and Llama 3.2 on the OpenWebText (Gokaslan2019, ) dataset. The dataset consists of 8,012,769 documents (9,035,582,198 tokens), out of which 4,007 documents are used for building the test dataset. All GPT2 models are trained on 4,915,200,000 tokens (4.9B) and Llama 3.2 is trained on 4,194,304,000 (4.2B) tokens. The test dataset is the same for all models. Section 5.3.2 shows the performance and accuracy for pretraining GPT2 and Llama 3.2 models. The parameters used for controlling the sparsity are found in the ablation studies in Section 5.4.
@@ 5.3.2. Accuracy and Performance of Pretraining
{tblr}
cells = c, cell21 = r=3, cell29 = r=3, cell210 = r=3, cell51 = r=2, cell59 = r=2, cell510 = r=2, hline1-2,5,7 = -, Model & $b$ $s_{max}$ $step\_size$ $d$ $L$ BLaST
BLaST
PPL
Dense
Dense
@@
Table 2. Pretraining results for BLaST with GPT2-XL and Llama 3.2 on OpenWebText. BLaST reduces end-to-end time of pretraining while preserving accuracy.
@@
Figure 8(a) shows the time per iteration over 4.9 billion tokens for the GPT2-XL model. The performance is compared for various sparsity configurations of BLaST against the dense case. The spikes observed for the sparse configurations are caused by mask generation, which occurs every 100 iterations. Dense matrix multiplication is used until the sparsity reaches 60%. At 60% sparsity, BSpMM routines (Section 3.3) become active, which explains the rapid decrease in per-iteration time as the sparsity increases. BLaST-80%/128x128 (plot ②) and BLaST-75%/128x128 (plot ③) configurations exhibit an earlier activation of SpMM routines due to higher values of $d$. Figure 8(b) shows the pretraining performance of Llama 3.2 over 4.2 billion tokens.
In Table 2, $s_{max}$ represents the maximum permissible sparsity in a given linear layer (see Equation 2). The sparsity-controlling hyperparameters include $b$ (the block size - Section 5.4.1), $step\_size$ (Section 5.4.2), $d$ (the sparsity decay - Section 5.4.3), and $L$ (the number of dense layers - Section 5.4.4).
5.3.3. Pretraining ViTs
In addition, we demonstrate that BLaST can be applied not only to NLP tasks but also to the vision domain. We train a ViT (dosovitskiy2020image, ) on CIFAR-10 (krizhevsky2009learning, ). Table 3 shows a comparison of four different ViT architectures under increasing sparsity levels achieved by BLaST.
Model Dense BLaST-70% BLaST-80% BLaST-90% BLaST-95%
ViT-B/16 98.57% 93.45% 93.82% 94.56% 93.08%
ViT-B/32 98.54% 92.91% 92.95% 92.70% 91.49%
ViT-L/16 98.88% 95.82% 95.71% 95.48% 95.31%
ViT-L/32 98.75% 92.13% 93.36% 93.59% 93.57%
Table 3. Accuracy of training four ViT architectures with BLaST. ViT is pretrained on a large vision task (dosovitskiy2020image, ), demonstrating BLaST’s knowledge transfer capability.
These results demonstrate the robustness of BLaST in handling increased sparsity levels in the MLP blocks of the ViT architecture.
Furthermore, in Figure 9 we illustrate how test accuracy evolves for both BLaST and the dense network as we iteratively train and sparsify the ViT architecture on CIFAR-10. We plot the test accuracy after every epoch (i.e., after each complete pass through the training data). We observe two key trends. First, as training progresses, BLaST requires fewer FLOP to process the entire training set. Second, BLaST achieves higher accuracy with fewer FLOP, resulting in a superior accuracy-to-FLOP ratio compared to the dense network.
Figure 9. Training BLaST-90%/128x128 for ViT-B/16 on CIFAR-10. BLaST requires fewer PFLOP to process the entire training set as we sparsify the network. BLaST achieves higher accuracy with fewer FLOP, resulting in a better accuracy-to-FLOP ratio.
@@ 5.3.3. Pretraining ViTs
The previous results demonstrate that BLaST can be used to speed up the pretraining of Transformer based architectures with negligible loss of accuracy.
5.4. Ablation Studies
5.4.1. Use of Block Sparsity
Table 4 shows the perplexity of the GPT2 models for different block sizes $b\times b$ (see Section 3.2). Every measurement is performed by capping the maximum sparsity $s_{max}$ of the dense matrices of the MLP layers at 70%. Increasing the value of $s_{max}$ obscures the effect of varying block sizes due to a deterioration in accuracy (Hoefler2021, ). The sparsity mask is updated at every iteration (i.e. $step\_size$ from Listing LABEL:lst:loop-prune-and-grow is set to 1).
Model Dense
BLaST
1x1
BLaST
16x16
BLaST
32x32
BLaST
64x64
BLaST
128x128
GPT2-small 6.48 9.30 7.53 7.51 7.49 7.44
GPT2-medium 5.54 7.26 6.04 6.02 6.00 6.04
GPT2-large 5.02 6.43 5.47 5.42 5.49 5.45
GPT2-XL 4.79 5.95 5.15 5.11 5.13 5.13
Table 4. Pretraining perplexity of the test data set for various block sizes compared to dense with maximum sparsity $s_{max}$ at 70% for all BLaST measurements.
All models exhibit improved perplexity with blocked sparsity. This improvement can be attributed to the ”Difference” step shown in Figure 2. During the regrowth step in Figure 2, the regrown blocks alter the final sparsity of the new weight matrix $W_{new}$.
Figure 10 presents the ratio of regrown blocks to the total number of blocks for block sizes $b$ ranging from $1$ to $128$, based on GPT2-small measurements from Table 4. When $b=1$, the ratio attains its maximum value and exhibits the greatest variation. In contrast, the blocked variants with $b$ values from $16$ to $128$ show more stable and lower ratios, which correlates with improved perplexity.
The proportion of regrown blocks serves as an indicator of whether the pruning of the weight matrices introduces significant discontinuities in the optimization function (Hu2024, ). A lower proportion of regrown blocks suggests that the weight pruning is more consistent with the gradient’s optimization direction.
Figure 10. Proportion of regrown blocks for $b\in\{1,16,32,64,128\}$ on GPT2-small.
5.4.2. Sparsity Step Size
The $step\_size$ variable in Listing LABEL:lst:loop-prune-and-grow controls the number of iterations after which the sparse masks are generated using the blocked prune and grow algorithm from Section 3.2. The pruning function $S()$ in Figure 2 introduces a discontinuity in the optimization of the neural network (Hu2024a, ). The larger the interval between the mask generation steps, the more discontinuous the pruning function (Hu2024a, ). However, generating sparse masks introduces a non-trivial overhead into the overall computation. Table 5.4.2 tests the perplexity for $step\_size$ varying from 1 to 1,000.
{tblr}
colspec = c *8c, booktabs, cells = c $step\_size$ 1 2 5 10 25 50 100 1000
GPT2-small 7.51 7.52 7.51 7.50 7.53 7.54 7.50 9.83
GPT2-medium 6.02 6.02 6.03 6.03 6.03 6.01 6.03 6.51
GPT2-large 5.42 5.43 5.42 5.42 5.42 5.41 5.42 5.70
GPT2-XL 5.12 5.11 5.11 5.11 5.12 5.11 5.12 5.26
Table 5. Pre-training perplexity of BLaST versus pruning interval ($step\_size$) with fixed $32\times 32$ blocks. Results show robustness up to $step\_size=100$.
The perplexity is almost the same for a given model for all values of $step\_size$ except at 1,000 iterations. This experimentally proves that the overhead of generating sparse masks can be reduced without losing accuracy by using larger values of $step\_size$.
5.4.3. Changing the Sparsity Decay
Modifying $d$ in Equation 2 triggers the use of BLaST BSpMM routine earlier in training, speeding up the entire pretraining process. Table 6 illustrates the impact of varying $d$ for the GPT2-large model.
{tblr}
colspec = *6c, booktabs, cells = c Decay parameter $d$ 0 1000 4000 7000 9000
Perplexity 5.42 5.44 5.44 5.44 5.43
Table 6. BLaST perplexity for GPT2-large when changing $d$ from Equation 2 after training with 4.9 billion tokens.
The models in Table 6 are trained on 4.9 billion tokens, which corresponds to $m=10,000$. For the GPT2 models, the block size $b$ is set to $32\times 32$. Note that the drop in accuracy is negligible compared to the case of $d=0$, as shown in Table 4. This indicates that the BLaST BSpMM can be applied earlier in the pretraining process without sacrificing accuracy compared to $d=0$.
5.4.4. Effect of Dense Layers
Previous work (nvidia2025prune, ) has shown that pruning the first and the last dense layers has the greatest impact on the model’s performance. In Figure 11, we retain both the first and the last dense layers while applying sparsification to either the leftmost or the rightmost MLP blocks of the GPT2-small model. These results empirically indicate that the improvement in perplexity is greater when the dense MLP blocks are positioned on the right. Consequently, we employ dense layers on the right side of the model to enhance accuracy in the performance runs described in Section 5.3.2. The number of dense layers on the right side is denoted by $L$ in Table 2.
Figure 11. Impact on the perplexity when keeping a certain number of dense MLP blocks on the left, and right side. This ablation study determines the $L$ hyperparameter for pretraining in Section 5.3.
@@ 6. Related Work
Sparse matrix multiplication on parallel architectures remains an active research area. Surveys summarize many effective techniques (gao2023, ; FilipponeReview, ), while recent work exploits dense MMA units to accelerate sparse computations (lu2023, ; niu2020, ; Zachariadis2020, ; Okanovic2024, ). For instance, DASP (lu2023, ) employs MMA units for SpMV acceleration, further boosted by CluB, which exposes dense substructures in sparse matrices. Reordering methods are key to improving data locality, bandwidth, and load balancing in sparse kernels (peng2020, ; Pichel2005, ; Pinar99, ; spMMHong, ; gianinazzi2024arrow, ). Graph-partitioning has boosted SpMV on multicore and GPU platforms (Trotter2023, ; Pichel2012, ; pichon2017, ; silva2017influence, ) using techniques such as Reverse Cuthill–McKee (RMC, ), approximate minimum degree (Amestoy2004, ), nested dissection (karypis1998fast, ), and Gray code ordering (Zhao2020, ). Recent advances, including community-based and just-in-time parallel reordering (Rabbit Order) (Balaji2023, ; rabbit, ) and others (gleinig2022optimal, ), further enhance cache locality and performance. Similarity-based reordering has been integrated into novel SpMM kernel designs. Hierarchical clustering via locality-sensitive hashing supports adaptive sparse tiling (peng2020, ; spMMHong, ), while similarity-based clustering improves both variable-size block multiplication (sylos2022, ), and the use of matrix units in block-based SpMM kernels (Okanovic2024, ).
@@ 7. Conclusion
This paper introduces BLaST, a general, scalable, and widely applicable end-to-end method that greatly reduces memory footprint, hardware cost of running contemporary extreme-scale LLMs, while simultaneously reducing runtime in all scenarios, both training- and inference-related. BLaST iteratively sparsifies the weights of the MLP layers inside the Transformer blocks, achieving up to 95% sparsity. Additionally, we introduce a highly efficient SpMM kernel that can be used as a standalone sparse linear algebra primitive, beyond ML workloads. We perform a comprehensive performance study of the proposed sparse kernel on state-of-the-art neural network architectures. We show a speedup of matrix multiplication of up to 16.7x, and we also demonstrate a speedup in the MLP block, achieving up to 8.8x improvement on the Llama-3.1 405B model. For Llama-3.2 1B, we show a 1.6x speedup in inference compared to the dense model, as well as a 1.1x speedup in pretraining. BLaST leads to memory savings of up to 3.12x during inference, translating into 2.9x lesser investment on GPUs and lower operating costs for data centers. We evaluate BLaST on 9 Transformer models across 8 datasets in both the text and vision domains, showing negligible accuracy losses compared to dense models. The presented results confirm that BLaST can be used as a drop-in method to scale up, and improve the efficiency of any Transformer-based architecture.

### v2

@@ 5.4.2. Text Classification (GLUE)
The GLUE benchmark (wang2019gluemultitaskbenchmarkanalysis, ) assesses the general language understanding capabilities of models. Specifically, we evaluate on five task: Recognizing Textual Entailment (RTE), the Multi-Genre Natural Language Inference Corpus (MNLI), The Stanford Sentiment Treebank (SST-2), the Winograd Schema Challenge (WNLI), and report the average score calculated as the arithmetic mean of all other task specific metrics.
In Table 2 we compare the performance of the Llama 3.2-1B model when sparsified with the P&G algorithm (Section 3.2) under different levels of sparsity, and various block sizes. We observe that P&G is robust to variations in block size, achieving a similar average score across the five tasks regardless of the chosen block size. Furthermore, the sparse models maintain accuracy comparable to that of the dense baseline. For instance, on the WNLI dataset, P&G achieves performance on par with the dense model across all settings. In all experiments, we iteratively sparsify the model during fine tuning to recover the accuracy of the dense version (see Section 3).
P&G (3.2) SST-2 MRPC RTE WNLI Avg.
$s_{max}$ $b$ ACC ACC / F1 ACC ACC Score
70% 32 84.98 72.06/80.94 59.57 56.34 69.97
80% 32 85.89 68.14/77.35 61.01 56.34 68.90
90% 32 84.63 70.34/79.17 61.37 56.34 69.19
95% 32 85.32 70.83/79.73 59.21 56.34 69.13
70% 64 84.40 66.42/77.28 59.21 56.34 67.77
80% 64 85.67 69.36/79.13 60.29 56.34 69.12
90% 64 84.29 68.87/77.36 58.12 56.34 67.75
95% 64 85.32 68.38/78.96 55.23 56.34 67.78
70% 128 86.24 70.83/79.86 60.65 56.34 69.84
80% 128 85.55 70.59/79.87 53.07 56.34 68.13
90% 128 85.09 69.12/80.25 59.21 56.34 69.02
95% 128 83.94 68.87/79.81 54.87 56.34 67.72
Dense 87.04 73.28/81.22 62.45 56.34 71.08
Table 2. Fine tuning Llama 3.2 1B with blocked prune and grow for GLUE benchmark. Blocked prune and grow is robust to the choice of tile size, capable of lossless sparsification up to 95% for WNLI.
Table 3 presents the accuracy results for the BERT-335M(Devlin2019, ) model on a subset of the GLUE benchmark using oLLM (Section 3.3). Although the average score for Llama 3.2-1B in Table 2 using P&G falls below the dense baseline, the sparsified BERT model occasionally surpasses dense performance. As observed in the text summarization task from Section 5.4.1, oLLM’s improved curvature estimation yields better accuracy under blocked sparsity.
@@
Collectively, these results demonstrate the robustness of the proposed sparsification algorithms to varying sparsity levels, and block sizes for reconstructing accuracy, making BLaST a ready to use method for pretrained networks when speedup and memory reduction are of importance.
oLLM (3.3) SST-2 MRPC RTE WNLI Avg.
$s_{max}$ $b$ ACC ACC / F1 ACC ACC Score
70% 1 90.71 68.38/81.22 55.95 56.33 67.66
80% 1 87.72 68.38/81.24 55.95 56.33 66.66
90% 1 89.44 68.38/81.29 53.79 56.33 66.52
95% 1 88 68.38/81.34 53.06 56.33 65.79
70% 16 88.64 68.38/81.22 53.79 56.33 66.25
80% 16 87.38 68.38/81.25 55.95 56.33 66.55
90% 16 83.25 68.38/81.30 53.42 56.33 64.33
95% 16 84.74 68.38/81.35 53.42 56.33 64.83
70% 32 89.79 68.38/81.22 53.42 56.33 66.51
80% 32 84.28 68.38/81.26 55.59 56.33 65.4
90% 32 83.37 68.38/81.31 53.79 56.33 64.49
95% 32 85.8 68.38/81.36 54.15 56.33 65.42
70% 64 89.9 68.38/81.22 53.79 56.33 66.67
80% 64 85.34 68.38/81.27 56.67 56.33 66.11
90% 64 88.76 68.38/81.32 55.95 56.33 67.01
95% 64 87.5 68.38/81.37 53.06 56.33 65.63
70% 128 90.82 68.38/81.23 53.79 56.33 66.98
80% 128 86.92 68.38/81.28 54.15 56.33 65.8
90% 128 87.72 68.38/81.33 53.06 56.33 65.70
95% 128 85.73 68.38/81.38 54.87 56.33 65.64
Dense 92.88 68.38/81.39 47.5 56.33 65.57
Table 3. Performance of BERT-335M on GLUE when fine-tuned with the oLLM pruning method (Section 3.3). Across a broad range of block sizes, oLLM maintains accuracy close to, or above the dense model, illustrating that second-order, block-structured pruning can be applied without substantial degradation on standard language understanding tasks.
5.4.3. Image Recognition (CIFAR10)
We demonstrate that P&G (Section 3.2) can be applied not only to NLP tasks but also to the vision domain. We fine-tune a pretrained ViT (dosovitskiy2020image, ) on CIFAR-10 (krizhevsky2009learning, ), using weights originally trained on ImageNet (deng2009imagenet, ). Table 4 shows a comparison of four different ViT architectures under increasing sparsity levels achieved by P&G. These results demonstrate the robustness of BLaST in handling increased sparsity levels in the MLP blocks of the ViT architecture.
Model Dense P&G-70% P&G-80% P&G-90% P&G-95%
ViT-B/16 98.57% 93.45% 93.82% 94.56% 93.08%
ViT-B/32 98.54% 92.91% 92.95% 92.70% 91.49%
ViT-L/16 98.88% 95.82% 95.71% 95.48% 95.31%
ViT-L/32 98.75% 92.13% 93.36% 93.59% 93.57%
Table 4. Accuracy of training four ViT architectures with P&G (3.2). ViT is pretrained on a large vision task (dosovitskiy2020image, ), demonstrating the knowledge transfer capability of this sparsification algorithm.
Furthermore, in Figure 10 we illustrate how test accuracy evolves for P&G and the dense network as we iteratively train and sparsify the ViT architecture on CIFAR-10. We plot the test accuracy after every epoch (i.e., after each complete pass through the training data). We observe two key trends. First, as training progresses, P&G requires fewer FLOP to process the entire training set. Second, P&G achieves higher accuracy with fewer FLOP, resulting in a superior accuracy-to-FLOP ratio compared to the dense network.
Figure 10. Training 90%/128x128 for ViT-B/16 on CIFAR-10 with P&G (3.2). Fewer PFLOP to process the entire training set are required with sparsification as we sparsify the network. P&G achieves higher accuracy with fewer FLOP, resulting in a better accuracy-to-FLOP ratio.
5.5. Memory Footprint for Inference
@@
Pruning weights by up to 95% significantly reduces the total parameter count and correspondingly decreases the model’s memory footprint. Figure 11 shows the number of NVIDIA GH200 GPUs required to store model weights, assuming 96GB HBM per GPU and FP32 precision. BLaST reduces the GPU requirements to approximately one-third of those needed for the dense Llama 3.2 405B model. This reduction directly translates to lower capital investments for inference data centers and decreased operational energy consumption.
@@
Figure 11. Number of GH200 GPUs required for storing weights in FP32 of the given model assuming 96GB per GPU. BLaST leads up to 4.45x inference memory reduction footprint.
@@ 5.6. BLaST for Pretraining
Figure 12. GPT2-XL pretraining with 4.9B tokens.
Figure 13. Time per iteration for GPT2-XL with OpenWebText (Gokaslan2019, ). Each line plot uses the hyper parameters from Table 5.
5.6.1. Causal Language Modelling
We pretrain GPT2-XL on the OpenWebText (Gokaslan2019, ) dataset with Causal Language modelling (CLM). The dataset consists of 8,012,769 documents (9,035,582,198 tokens), out of which 4,007 documents are used for building the test dataset. The model is trained on 4,915,200,000 tokens ($\approx$4.9B). Section 5.6.2 shows the performance and accuracy for pretraining GPT2 and Llama 3.2 models. We use only the P&G algorithm for pretraining due to the higher computational cost of oLLM, as shown in Section 3.3.
@@ 5.6.2. Accuracy and Performance of Pretraining
Model $b$ $s_{max}$ step_sizestep\_size $d$ $L$ BLaST
(h) BLaST
PPL Dense
(h) Dense
@@
Table 5. Pretraining results for P&G (3.2) with GPT2-XL on OpenWebText. BLaST reduces end-to-end time of while preserving accuracy.
@@
Figure 13 shows the time per iteration over 4.9 billion tokens for the GPT2-XL model. The performance is compared for various sparsity configurations of BLaST against the dense case. The spikes observed for the sparse configurations are caused by mask generation, which occurs every 100 iterations. Dense matrix multiplication is used until the sparsity reaches 60%. At 60% sparsity, the BSpMM routines (Section 3.4) take over, which leads to a sharp drop in per-iteration time as sparsity increases. BLaST-80%/128x128 (plot ②) and BLaST-75%/128x128 (plot ③) configurations exhibit an earlier activation of SpMM routines due to higher values of $d$.
@@
The previous results demonstrate that BLaST with P&G can be used to speed up the pretraining of Transformer based architectures with negligible loss of accuracy.
@@ 6. Related Work
Sparse matrix multiplication on parallel architectures remains an active research area. Surveys summarize many effective techniques (gao2023, ; FilipponeReview, ), while recent work exploits dense MMA units to accelerate sparse computations (lu2023, ; niu2020, ; Zachariadis2020, ; Okanovic2024, ). For instance, DASP (lu2023, ) employs MMA units for SpMV acceleration, further boosted by CluB, which exposes dense substructures in sparse matrices. Reordering methods are key to improving data locality, bandwidth, and load balancing in sparse kernels (peng2020, ; Pichel2005, ; Pinar99, ; spMMHong, ; gianinazzi2024arrow, ; Ma2025b, ). Graph-partitioning has boosted SpMV on multicore and GPU platforms (Trotter2023, ; Pichel2012, ; pichon2017, ; silva2017influence, ) using techniques such as Reverse Cuthill–McKee (RMC, ), approximate minimum degree (Amestoy2004, ), nested dissection (karypis1998fast, ), and Gray code ordering (Zhao2020, ). Recent advances, including community-based and just-in-time parallel reordering (Rabbit Order) (Balaji2023, ; rabbit, ) and others (gleinig2022optimal, ), further enhance cache locality and performance. Similarity-based reordering has been integrated into novel SpMM kernel designs. Hierarchical clustering via locality-sensitive hashing supports adaptive sparse tiling (peng2020, ; spMMHong, ), while similarity-based clustering improves both variable-size block multiplication (sylos2022, ), and the use of matrix units in block-based SpMM kernels (Okanovic2024, ).
@@ 7. Conclusion
This paper introduces BLaST, a general, scalable, and widely applicable end-to-end framework that greatly reduces memory footprint, hardware cost of running contemporary large-scale LLMs, while simultaneously reducing runtime in all scenarios, both training- and inference-related. BLaST iteratively sparsifies the weights of the linear layers inside the Transformer blocks, achieving up to 95% sparsity with negligible accuracy loss. Beyond its current implementation, BLaST provides a flexible foundation that can seamlessly integrate with any future block pruning algorithm, ensuring its continued relevance as sparsification techniques evolve. BLaST introduces a highly efficient BSpMM kernel that can be used as a standalone sparse linear algebra primitive, beyond ML workloads. We perform a comprehensive performance study of the proposed sparse kernel on state-of-the-art neural network architectures. The kernel delivers up to $521\times$ speedup over state of the art SpMM, up to $23\times$ over dense GEMM, $1.6\times$ end to end speedups, and $2.2\times$ distributed inference speedups on the Llama3 family. These gains hold across a range of block sizes, which provide a practical knob to trade speed for accuracy. We see BLaST as a reliable building block for efficient inference and training pipelines at large scale.